OpenStudy (anonymous):

A truck with 0.420-m-radius tires travels at 32.0 m/s. What is the angular velocity of the rotating tires in radians per second? What is this in rev/min?

OpenStudy (jango_in_dtown):

\[v=r \omega\] where v is velocity and r is radius and w is angular velocity

OpenStudy (jango_in_dtown):

here v=32m/s r=0.420m so w=v/r=32/0.42=76.2 rad/sec (approx.) now 1rev=2pi rad hence in rev/sec we have 76.2rad/s *(1rev/2pi rad)*(60s/1min)=728rev/s

OpenStudy (anonymous):

I got it. Thanks for your help, jango_IN_DTOWN

OpenStudy (jango_in_dtown):

no problem...:)